# REAR: A Relevance-Aware Retrieval-Augmented Framework for Open-Domain Question Answering

### v1

@@ Abstract
Considering the limited internal parametric knowledge, retrieval-augmented generation (RAG) has been widely used to extend the knowledge scope of large language models (LLMs). Despite the extensive efforts on RAG research, in existing methods, LLMs cannot precisely assess the relevance of retrieved documents, thus likely leading to misleading or even incorrect utilization of external knowledge (i.e., retrieved documents). To address this issue, in this paper, we propose REAR, a RElevance-Aware Retrieval-augmented approach for open-domain question answering (QA). As the key motivation, we aim to enhance the self-awareness of source relevance for LLMs, so as to adaptively utilize external knowledge in RAG systems. Specially, we develop a new architecture for LLM based RAG system, by incorporating a specially designed rank head that precisely assesses the relevance of retrieved documents. Furthermore, we propose an improved training method based on bi-granularity relevance fusion and noise-resistant training. By combining the improvements in both architecture and training, our proposed REAR can better utilize external knowledge by effectively perceiving the relevance of retrieved documents. Experiments on four open-domain QA tasks show that REAR significantly outperforms previous a number of competitive RAG approaches. Our code and data can be accessed at https://github.com/RUCAIBox/REAR.
@@ 1 Introduction
Despite the excellent capacities, large language models (LLMs) Brown et al. (2020); Zhao et al. (2023) still struggle with knowledge-intensive tasks like open-domain question answering (QA), lacking in real-time and domain knowledge Li et al. (2023a). To mitigate this issue, retrieval-augmented generation (RAG) provides LLMs with potentially relevant documents through retrieval Gao et al. (2023), aiding in generating more precise content.
@@
While RAG offers clear benefits, it also introduces several technical challenges for effectively improving LLMs. Firstly, the retrieved results likely contain irrelevant content or documents, which may mislead LLMs and even cause them to respond incorrectly Ren et al. (2023); Mallen et al. (2023). Moreover, it has become common to incorporate multiple reference documents to boost the overall reliability of retrieved documents, whereas this approach potentially amplifies the impact of the noise present in the retrieved documents Liu et al. (2023); Shi et al. (2023). Thus, LLMs face difficulties in filtering irrelevant documents and integrating their internal knowledge, which needs to avoid potential interference with noisy content.
@@
Recently, several studies Luo et al. (2023); Asai et al. (2023); Yoran et al. (2023) have attempted to enhance the robustness of RAG systems. For instance, Self-RAG Asai et al. (2023) allows the model to introspect its outputs by generating special tokens to discriminate if the documents are relevant, and RobustLM Yoran et al. (2023) prompts LLMs to first discriminate if the documents are relevant and then generate answers. However, these approaches perform the assessment of document relevance solely based on binary labels, which are highly sparse and not precise to capture the fine-grained relevance. In addition, they seldom consider the varied relevance degree of reference documents, making the utilization somehow blind.
@@
To this end, in this paper, we propose REAR, a RElevance-Aware Retrieval-augmented approach for open-domain question answering (QA). The key idea of our approach is to develop strong self-awareness of source relevance for RAG systems, so that the LLM can learn to adaptively utilize external knowledge (i.e., retrieved documents) for solving complex QA tasks. To achieve this goal, we make two major contributions in both model architecture and training. Firstly, inspired by the success of reranker in the field of information retrieval, which can discern the differences in relevance among retrieved documents with fine-grained preference Zhao et al. (2023), we fuse LLMs with a specially designed rank head for relevance assessment, and then employ it to capture relevance signals, to avoid distractions from irrelevant external knowledge. Secondly, to overcome the limitations of binary discriminative methods that can only provide coarse-grained supervision signals, we further design bi-granularity relevance fusion by additionally leveraging fine-grained supervision, and also enhance the discrimination ability of RAG systems by noise-resistant training.
To summarize, our proposed framework enhances the self-awareness of source relevance when using reference documents in RAG systems, with technical improvements on both model architecture and training. Extensive experiments across four different open-domain QA tasks demonstrate the effectiveness of our REAR framework.
@@ 3 Task Formulation
Formally, given a query $q$, the retriever outputs top-$k$ documents $\mathcal{D}=\{d_{i}\}_{i=1}^{k}$ from a document collection (can be refined by an optional reranker) at the first stage. Different from prior work that combines the entire set of retrieved documents as a unified reference for answer generation Luo et al. (2023); Xu et al. (2023); Hofstätter et al. (2023), our approach emphasizes individual document utilization, which can be also extended to a multi-document setting. Given the input query $q$ and reference documents $\mathcal{D}=\{d_{i}\}_{i=1}^{k}$, the reader (i.e., the LLM) generates an answer $a_{i}$ based on each reference document $d_{i}$, forming an answer set $\mathcal{A}$:
@@ 4 Methodology
In this section, we present the proposed RElevance-Aware Retrieval-augmented generation framework (REAR), which is capable of precisely assessing the relevance between queries and documents by incorporating a reranker module within the LLM. Furthermore, we propose improved training methods to enhance the performance of the entire framework, including bi-granularity relevance fusion and noise-resistant training.
@@ 4.1 Relevance-Aware RAG Architecture
In this part, we propose a novel architecture that augments the LLM with a ranking module for enhancing the awareness of relevance. As shown in Figure 1, the architecture encompasses three parts, including relevance assessment, relevance-guided generation, and final answer routing.
@@ 4.1.1 Relevance Assessment
Instead of treating the retrieved documents equally, we first aim to assess the relevance scores of query-document pairs with the incorporation of a reranker module. Drawing from the success of rerankers in achieving precise relevance assessment Sun et al. (2023); Ma et al. (2023), we introduce a rank head upon the LLM, designed for capturing relevance signals through full interactions between queries and documents. For assessing the relevance score, the LLM decoder maps the input query-document pair into an embedding $\bm{v}_{\text{rel}}$:
$\bm{v}_{\text{rel}}=\text{Decoder}(q,d).$ (2)
Subsequently, $\bm{v}_{\text{rel}}$ is quantified into a score $s_{\text{rel}}$ by a specially designed rank head:
$s_{\text{rel}}=\text{RankHead}(\bm{v}_{\text{rel}}),$ (3)
where $\text{RankHead}(\cdot)$ is implemented as a linear projection layer. This design offers two advantages. Firstly, the rank head specializes in relevance assessment and does not affect the functionalities of other modules in the LLM. Secondly, the relevance score $s_{\text{rel}}$ can be directly optimized based on the internal states of the LLM with various loss functions, such as contrastive learning and preference optimization, making the training of relevance assessment more flexible and concentrated (will be introduced in Section 4.2).
@@ 4.1.2 Relevance-guided Generation
Different from previous work that ignores the relevance of document Cuconasu et al. (2024), we aim to integrate the relevance assessment score of each document into LLMs to guide the answer generation process. Since the relevance score $s_{\text{rel}}$ (in Eq. 3) is a scalar, which may not be fully utilized by LLMs, we further incorporate an embedding layer to map it into a dense vector $\bm{v}_{\text{guide}}$ as:
@@
Note that we do not directly utilize the dense vector $\bm{v}_{\text{rel}}$ (Eq. 2) as the signal $\bm{v}_{\text{guide}}$ to guide the answer generation (Eq. 5). The reasons are twofold. First, $\bm{v}_{\text{rel}}$ plays the role of query-document relevance assessment, while $\bm{v}_{\text{guide}}$ is supposed to guide LLMs to generate correct answers. Thus, employing different relevance representations $\bm{v}_{\text{rel}}$ and guide signal $\bm{v}_{\text{guide}}$ can avoid the interference between relevance assessment and answer generation. Second, our experiments show that utilizing distinct $\bm{v}_{\text{rel}}$ and $\bm{v}_{\text{guide}}$ in our framework leads to a superior performance (see Section 5.3.1).
4.1.3 Final Answer Routing
After the retrieval process, we can obtain the top-$k$ relevant documents $\mathcal{D}=\{d_{i}\}_{i=1}^{k}$. Based on Eq. 2-Eq. 5, we conduct inference on each query-document pair $(q,d_{i})$ to obtain corresponding answer $a_{i}$, ended up with a set of $k$ answers $\mathcal{A}=\{a_{i}\}_{i=1}^{k}$. We view the generation process from different retrieved documents as navigating distinct reasoning paths, and we aim to select the most reliable routing path to obtain the final answer. Specially, we propose two routing strategies, namely path-reliability and knowledge-consistency.
• Path-reliability routing: A straightforward approach is to select the answer that the LLM deems to have the highest relevance score:
$a^{*}=\operatorname*{arg\,max}_{a_{i}\in\mathcal{A}}s_{\text{rel}}(q,d_{i}),$ (6)
where $s_{\text{rel}}(q,d_{i})$ represents the relevance score of pair $(q,d_{i})$ obtained by Eq. 3.
• Knowledge-consistency routing: Inspired by the success of self-consistency in Chain-of-Thought reasoning Wei et al. (2022); Wang et al. (2022), this approach aims to select the answer that is based on the external knowledge (i.e., retrieved documents), which is highly consistent with the LLMs’ internal knowledge (i.e., parametric knowledge). In particular, we set the relevance score to zero (denoted by $\hat{s}_{\text{rel}}$) and calculate the inverse of perplexity $c_{i}$ Meister and Cotterell (2021) of generating the answer $a_{i}$ as follows:
$c_{i}=\frac{1}{\text{PPL}(a_{i}\mid q,d_{i},\hat{s}_{\text{rel}}=0)}.$ (7)
By setting the relevance score to zero, the LLM tends to generate the answer $a_{i}$ based on its internal knowledge instead of the retrieved document $d_{i}$. If the answer $a_{i}$ can also be generated with lower perplexity (higher knowledge-consistency score $c_{i}$) when $\hat{s}_{\text{rel}}=0$, it indicates that the external knowledge is highly consistent with the internal knowledge. Thus, we first select a subset of documents with high relevance by a threshold $\gamma$:
$\mathcal{D}^{+}=\{d_{i}\mid s_{\text{rel}}(q,d_{i})>\gamma\}.$ (8)
Then, we linearly combine the knowledge-consistency score $c_{i}$ with the relevance score $s_{\text{rel}}(q,d_{i})$ with a hyperparameter $\lambda$ to select the final answer, denoted as $a_{c}^{*}$:
$a^{*}_{\text{c}}=\operatorname*{arg\,max}_{\{a_{i}\mid d_{i}\in\mathcal{D}^{+}\}}(\text{~{}}s_{\text{rel}}(q,d_{i})+\lambda c_{i}).$ (9)
@@ 4.2 Model Training
In this part, we will introduce the training method for optimizing our approach.
@@ 4.2.1 Bi-granularity Relevance Fusion
Precise relevance assessment is crucial for the reliable utilization of retrieved documents. Previous work often adopts the coarse-grained binary discrimination task Yoran et al. (2023), which cannot provide sufficient evidence for solving complex QA tasks. Therefore, we consider further incorporating a fine-grained ranking optimization objective, to improve the model training.
Specially, for the coarse-grained supervision, we set the binary labels for candidate documents $\mathcal{D}=\{d_{i}\}_{i=1}^{k}$, denoting “irrelevant” and “relevant” as $y\in\{0,1\}$, and optimize the following loss:
$\mathcal{L}_{\text{coarse}}=-\sum_{i=1}^{k}y_{i}\log\left(\sigma_{i}\right)+(1-y_{i})\log\left(1-\sigma_{i}\right),$ (10)
where $\sigma_{i}$ denotes the normalized probability of assessing $(q,d_{i})$ as relevant by the LLM.
For the fine-grained supervision, we utilize the estimated relevance scores $s_{\text{rel}}$ for deriving ranking preference constraints:
$\mathcal{L_{\text{fine}}}=-\sum_{i=1}^{k-1}\sum_{j=i+1}^{k}\log\left(\sigma_{i}-\sigma_{j}\right).$ (11)
Note that we assume that the complete annotation of partial preference relations for the retrieved documents $\mathcal{D}=\{d_{i}\}_{i=1}^{k}$ is available during training (See Section 4.2.3), where a smaller index indicates a more relevant document. Furthermore, we combine these two loss functions, as the objective of the bi-granularity relevance fusion:
@@ 4.2.2 Noise-resistant Training
$\mathcal{L}_{\text{noise-resistant}}=\sum_{d_{i}\in\mathcal{D}\cup\mathcal{D}^{-}}\log P(a\mid q,d_{i},s_{\text{rel}}).$ (13)
Through noise-resistant training, the LLM can learn to discern the incorporation of irrelevant documents, without being encumbered by extraneous information. It enforces LLMs to judge the relevance of external evidence, so as to decide whether to use the retrieved documents for answer generation.
@@
4.2.3 Training Data Construction
To optimize our model, we need high-quality training data (both positive and negative samples), and next introduce the concrete strategy for training data construction.
Relevance Labels Acquisition. To obtain fine-grained relevance labels used in Section 4.2.1, we employ a small-scale reranker with cross-encoder architecture to acquire the relevance assessments. In the field of information retrieval, the cross-encoder architecture is regarded as effective for assessing relevance degree, as it considers the query-document pair as a single text, and processes it via an encoder to obtain relevance scores $s_{\text{ce}}$ through full interactions Zhao et al. (2024); Santhanam et al. (2022). In combination with the traditional method of binary annotating label $y$, the generated score is given as:
$s_{\text{rel}}=\frac{1}{2}\left(s_{\text{ce}}+y\right),$ (15)
where $y\in\{0,1\}$ is the binary relevance label and $s_{\text{ce}}\in[0,1]$ is the continuous scores assessed by the aforementioned cross-encoder.
Irrelevant Documents Sampling. The training method necessitates the use of irrelevant (negative) documents. It has been shown that negative sampling has a large impact on relevance assessment Xiong et al. (2020). Specially, we refine SimANS Zhou et al. (2022) that ensures negatives are neither too difficult (false negatives) nor too trivial (uninformative):
where the sampling probability for the hard negative document is $p_{i}$, $s_{i}$ and $\hat{s}^{+}$ respectively denote the relevance scores of document $d_{i}$ and the positive document, and $a$, $b$, and $k$ are hyperparameters. As the relevance scores are evaluated by rerankers, the likelihood of high-scoring samples being false negatives increases. By incorporating a decay scaler $k$ into the sampling probability when relevance scores are high, we reduce the chance of sampling false negatives.
4.3 Model Discussion
Our primary contribution lies in the integration of a special ranking module (i.e., rank head in Section 4.1.1) with LLMs. In this way, relevance assessment and answer generation can be approached by the rank head and generation head, respectively, which can reduce the interference between the two different objects. Thus, this method more effectively assesses the fine-grained query-document relevance and further guides the answer generation.
Another key improvement is our training strategy. The bi-granularity relevance fusion enables the LLM to identify not just the overall relevance degree of documents but also to discriminate the relevance preference among the candidates. Consequently, our REAR shows superior performance in evaluating single documents and choosing among multiple documents. In noise-resistant generation training, the LLM is specifically taught to leverage relevance scores for the selective utilization of retrieved documents. This process fosters a dynamic capability to utilize both internal and external knowledge effectively.

### v2

@@ Abstract
Considering the limited internal parametric knowledge, retrieval-augmented generation (RAG) has been widely used to extend the knowledge scope of large language models (LLMs). Despite the extensive efforts on RAG research, in existing methods, LLMs cannot precisely assess the relevance of retrieved documents, thus likely leading to misleading or even incorrect utilization of external knowledge (i.e., retrieved documents). To address this issue, in this paper, we propose REAR, a RElevance-Aware Retrieval-augmented approach for open-domain question answering (QA). As the key motivation, we aim to enhance the self-awareness regarding the reliability of external knowledge for LLMs, so as to adaptively utilize external knowledge in RAG systems. Specially, we develop a novel architecture for LLM-based RAG systems, by incorporating a specially designed assessment module that precisely assesses the relevance of retrieved documents. Furthermore, we propose an improved training method based on bi-granularity relevance fusion and noise-resistant training. By combining the improvements in both architecture and training, our proposed REAR can better utilize external knowledge by effectively perceiving the relevance of retrieved documents. Experiments on four open-domain QA tasks show that REAR significantly outperforms previous a number of competitive RAG approaches. Our codes can be accessed at https://github.com/RUCAIBox/REAR.
@@ 1 Introduction
Despite the progressive capacities, large language models (LLMs) Brown et al. (2020); Zhao et al. (2023) still struggle with knowledge-intensive tasks like open-domain question answering (QA), lacking in real-time and domain knowledge Li et al. (2023a); Cheng et al. (2024). To mitigate this issue, retrieval-augmented generation (RAG) provides LLMs with potentially relevant documents through a retrieval module Gao et al. (2023), aiding in generating more precise content.
Figure 1: LLMs may be misled by irrelevant documents, and struggle to determine the relevance of a document Ren et al. (2023); Zhang et al. (2024).
@@
While RAG offers clear benefits, it also introduces several technical challenges for effectively improving LLMs. Firstly, the retrieved results likely contain irrelevant content or documents, which may mislead LLMs and even cause them to respond incorrectly Ren et al. (2023); Mallen et al. (2023). Moreover, it has become common to incorporate multiple reference documents to boost the overall reliability of retrieved documents. However, this approach potentially amplifies the impact of the noise present in the retrieved documents Liu et al. (2023); Shi et al. (2023). Thus, LLMs face difficulties in filtering irrelevant documents and integrating their internal knowledge Dong et al. (2023), which needs to avoid potential interference with noisy content.
@@
Recently, several studies Luo et al. (2023); Asai et al. (2023); Yoran et al. (2023) have attempted to enhance the robustness of RAG systems. For instance, Self-RAG Asai et al. (2023) allows the model to introspect its outputs by generating special tokens to discriminate if the documents are relevant, and RobustLM Yoran et al. (2023) prompts LLMs to first discriminate if the documents are relevant and then generate answers. However, these approaches perform the assessment of document relevance solely based on binary labels, which are highly sparse and not precise to capture the fine-grained relevance. In addition, they seldom consider the varied relevance degree of reference documents, making the utilization of external knowledge somehow blind.
@@
To this end, in this paper, we propose REAR, a RElevance-Aware Retrieval-augmented generation approach for open-domain question answering (QA). Our key idea is to develop robust self-awareness regarding the reliability of external knowledge (i.e., retrieved documents) within RAG systems, so that the LLM can learn to adaptively utilize the internal and external knowledge for solving complex QA tasks. To achieve this goal, we make two major contributions in both model architecture and training. First, we propose relevance-aware RAG architecture by incorporating explicit assessment modules in LLMs’ generation architecture to perform an additional relevance assessment task. In our architecture, the assessment module effectively captures relevance signals, and feeds them back to avoid distractions from irrelevant external knowledge during generation. Secondly, to support the relevance-aware RAG architecture, we further propose two training strategies. Bi-granularity relevance fusion strategy further integrates both coarse and fine-grained relevance supervision to overcome the limitations of binary discriminative methods, while noise-resistant training strategy enhances the discrimination ability of the LLM by incorporating negatives in the training procedure.
To the best of our knowledge, we are the first to introduce the idea of incorporating explicit assessment modules in the generation architecture of LLMs to aid in irrelevance-resistant generation. Extensive experiments on public open-domain QA benchmarks attest to the effectiveness of our REAR framework. Notably, we also demonstrate the strong generalization capability of REAR by conducting out-of-domain evaluation on multiple open-domain QA benchmarks.
@@ 3 Task Formulation
Formally, given a query $q$, the retriever outputs top-$k$ documents $\mathcal{D}=\{d_{i}\}_{i=1}^{k}$ from a document collection (can be refined by an optional reranker) at the first stage. Different from prior studies that combine the entire set of retrieved documents as a unified reference for answer generation Luo et al. (2023); Xu et al. (2023); Hofstätter et al. (2023), our approach emphasizes individual document utilization, which can be also extended to a multi-document setting. Given the input query $q$ and reference documents $\mathcal{D}=\{d_{i}\}_{i=1}^{k}$, the reader (i.e., the LLM) generates an answer $a_{i}$ based on each reference document $d_{i}$, forming an answer set $\mathcal{A}$:
@@ 4 Methodology
In this section, we present the proposed RElevance-Aware Retrieval-augmented generation framework (REAR), which is capable of precisely assessing the relevance degree during the generation process by incorporating explicit assessment modules within the LLM. Furthermore, we propose optimized training methods that are compatible with the REAR framework to support efficient operation, including bi-granularity relevance fusion and noise-resistant training.
@@ 4.1 Relevance-Aware RAG Architecture
In this part, we propose a novel architecture that augments the LLM with a relevance-assessing module for enhancing the awareness of irrelevant interference. As shown in Fig. 2 (a), the inference of the REAR architecture encompasses three steps, including relevance assessment, relevance-guided generation, and knowledge reliability verification.
@@ 4.1.1 Relevance Assessment
Instead of treating all the retrieved documents equally, we first aim to assess the relevance degrees of the documents. Drawing from the success of LLM-based decoder in achieving precise relevance assessment Sun et al. (2023); Ma et al. (2023), we first map the input query-document pair into the relevance embedding $\bm{v}_{\text{rel}}$ by the LLM:
$\bm{v}_{\text{rel}}=\text{LLM}(q,d)[-1].$ (2)
Subsequently, $\bm{v}_{\text{rel}}$ is quantified into a score $s_{\text{rel}}$ by the assessment module:
$s_{\text{rel}}=\text{Assess}(\bm{v}_{\text{rel}}),$ (3)
where $\text{Assess}(\cdot)$ is the assessment module implemented as a linear projection layer.
@@ 4.1.2 Relevance-guided Generation
Different from previous works that ignore the relevance of document Cuconasu et al. (2024), we aim to integrate the relevance score of each document into LLMs to assess document reliabilities and subsequently guide the generation process. Since the relevance score $s_{\text{rel}}$ (in Eq. 3) is a scalar, which may not be fully utilized by LLMs, we further incorporate an embedding layer to map it into a dense vector $\bm{v}_{\text{guide}}$ as:
@@
4.1.3 Knowledge Reliability Verification
Based on the generated answers, we finally verify the correctness of the answers by considering two factors: (a) Is the provided document reliable enough to trust the corresponding answer? (b) Without referring to the documents, to what degree will the LLM adhere to its original response? Specially, we propose two strategies, namely source-reliability and knowledge-consistency.
• Source-reliability: This strategy primarily emphasizes the quality of external knowledge. If an LLM assigns a high relevance score to a document, then the answer derived from it is considered more reliable.
• Knowledge-consistency : This approach further verifies if the provided knowledge conflicts with the parametric knowledge. Specifically, inspired by the success of self-consistency in Chain-of-Thought reasoning Wei et al. (2022); Wang et al. (2022), we inform the LLM that the document is irrelevant by setting the relevance score to zero (denoted by $\hat{s}_{\text{rel}}$) and calculate the inverse of perplexity $c$ Meister and Cotterell (2021) of generating the answer $a$:
$c=\frac{1}{\text{PPL}(a\mid q,d,\hat{s}_{\text{rel}}=0)},$ (6)
which evaluates the extent of LLM to stand by its original answer based on the parametric knowledge. Then, we linearly combine the knowledge-consistency score $c_{i}$ with the relevance score $s_{\text{rel}}(q,d_{i})$ to select the final answer.
@@ 4.2 Model Training
In this part, we will introduce the training pipeline for optimizing our approach, As shown in Fig. 2 (b).
@@ 4.2.1 Bi-granularity Relevance Fusion
Precise relevance assessment is crucial for the reliable utilization of retrieved documents. Previous work often adopts the coarse-grained binary discrimination task Yoran et al. (2023) , which cannot provide sufficient evidence for solving complex QA tasks. Therefore, we consider further incorporating a preference-based fine-grained task. Specifically, for the fine-grained supervision, we utilize the estimated relevance scores (See Section 4.2.3) for deriving relevance preference constraints:
$\mathcal{L_{\text{fine}}}=-\sum_{i}\sum_{j}(s_{i}>s_{j})\log\left(\sigma_{i}-\sigma_{j}\right),$ (7)
where $\sigma_{i}$ denotes the normalized probability of assessing $(q,d)$ as relevant by the LLM. Furthermore, we combine it with the coarse-grained binary loss $\mathcal{L_{\text{coarse}}}$, as the objective of the bi-granularity relevance fusion:
@@ 4.2.2 Noise-resistant Training
$\mathcal{L}_{\text{noise-resistant}}=\sum_{d\in\mathcal{D}\cup\mathcal{D}^{-}}\log P(a\mid q,d,s_{\text{rel}}).$ (9)
Through noise-resistant training, the LLM can learn to discern the incorporation of irrelevant documents, without being encumbered by extraneous information.
4.2.3 Training Data Construction
To optimize our model, we need high-quality training samples (both positive and negative samples) and labels.
Relevance Labels Acquisition. To obtain fine-grained relevance labels used in Section 4.2.1, we employ a small-scale reranker to acquire the continuous relevance score $s_{\text{ce}}$. We adopt rerankers with the cross-encoder architecture, since they are regarded as effective for assessing relevance degree Zhao et al. (2024); Khattab and Zaharia (2020). In combination with the traditional method of binary annotating label $y$, the estimated score is given as:
$s_{\text{rel}}=\frac{1}{2}\left(s_{\text{ce}}+y\right).$ (10)
This labeling approach combines lexical and semantic similarity, allowing for the acquisition of high-quality labels without accessing GPT APIs.
Irrelevant Documents Sampling. The training method necessitates the use of irrelevant (negative) documents. It has been shown that negative sampling has a large impact on relevance assessment Xiong et al. (2020). Specially, as shown in Fig. 2 (b), we refine SimANS Zhou et al. (2022) that ensures negatives are neither too difficult (false negatives) nor too trivial (uninformative):
where the sampling probability for the hard negative document is $p_{i}$, $s_{i}$ and $\hat{s}^{+}$ respectively denote the relevance scores of document $d_{i}$ and the positive document, and $a$, $b$, and $k$ are hyperparameters. By incorporating a decay scaler $k$ into the sampling probability when relevance scores are high, we reduce the chance of sampling false negatives.
@@ 4.2.3 Training Data Construction
4.3 Discussion
Aspect Self-RAG CoN SAIL REAR (ours)
Assess Gen Gen Gen Explicit Module
Train SFT SFT SFT SFT+ Contrastive Loss
Data GPT GPT GPT Free Model (110M)
Table 1: The difference between REAR and previous work. Assess, Train and Data are short for relevance assessment method, training loss, and data construction methods respectively. REAR utilizes an explicit module for relevance assessment, and adopts bi-granularity (involving contrastive loss) for training. Furthermore, we label the data without access to GPT APIs.
Distinctions from Existing Methods. As shown in Table 1, our primary contribution lies in the architecture design, which differs significantly from existing studies. Under our optimized architecture, LLMs can generate more fine-grained relevance signals to aid in the following generation process. Besides, LLMs can further calculate the consistency between parametric and external knowledge to evaluate the reliability of answers. Moreover, this architecture makes it easy to adopt the proposed preference-based and noise-resistant loss functions. Furthermore, our label machine makes good use of smaller models and traditional labels, and our sampling strategy improves training data quality, eliminating the need for GPT APIs. As a result, REAR achieves more precise relevance evaluation and better generation performance (Table 3).
Methods T.C. Training Inference
CoN $\mathcal{O}((p+nd)^{2})$ 10.34s/step 0.82s/sample
Self-RAG $\mathcal{O}(n(p+d)^{2})$ 6.52s/step 1.41s/sample
REAR (ours) $\mathcal{O}(n(p+d)^{2})$ 6.33s/step 0.45s/sample
Table 2: The efficiency analysis of REAR and previous work. T.C. is short for time complexity. $d$, $p$ and $n$ denote the length of the document, the length of the prompt, and the number of documents respectively.
Efficiency. We further discuss the efficiency of our REAR, as shown in Table 2. First, we compare REAR with other RAG frameworks that employ different task formulations, such as Chain-of-Note (CoN) Yu et al. (2023). CoN processes extensive paragraphs and generates in-depth analyses to identify usable parts of document collections. This methodology leads to increased training and inference times due to the quadratic time complexity associated with transformers Dong et al. (2024), where time is proportional to the square of the input sequence length. Besides, compared to Self-RAG, which follows a similar approach, REAR achieves a reduction in inference time. This improvement is primarily due to our integration of PagedAttention Kwon et al. (2023). By using PagedAttention, we ensure that calculations performed during the relevance assessment phase are preserved, thereby eliminating the need for redundant recalculations. The comparisons of actual training and inference times in Table 2 further illustrate the computational efficiency of our method.